OpenStudy (anonymous):

Find the equation of a parabola that passes through the points (-2, 16), (0, 2), and (1, 4)

OpenStudy (anonymous):

Let \[ f(x) = a x^2 + b x +c \] We need to find a, b, and c. We need three equations in three unknown a,b and c. Can you find these equations?

OpenStudy (anonymous):

f(-2) = 4 a - 2 b + c = 16 f(0) =c=2 f(1)= a+b+c =4

OpenStudy (anonymous):

You can see that c=2, so you can reduce the 3 equations to 2 equations. Now find a, b

OpenStudy (anonymous):

Finally c= 2, a = 3, b = -1

OpenStudy (anonymous):

Your parabola is \[ f(x) = 3 x^2 - x +2 \]

OpenStudy (anonymous):

Thanks